OpenStudy (anonymous):

Rationalize the denominator of 5/√3

OpenStudy (anonymous):

5 / sqrt 3 = (5 / sqrt 3)(sqrt 3 / sqrt 3) = 5 sqrt 3 / 3

OpenStudy (unklerhaukus):

\[\tt{Rationalising~~the~~denominator}\]_____________________ Any irrational fraction of the form \[\frac a{\sqrt b}\] can be rationalised by multiplying by \(\dfrac{\sqrt b}{\sqrt b}\) (which equals one) like this \[\large\frac a{\sqrt b}=\frac a{\sqrt b}\times1\\ \large\qquad=\frac a{\sqrt b}\times\frac{\sqrt b}{\sqrt b}\\ \large\qquad=\frac{a\sqrt b}{b}\] _____________________ If the irrational fraction of this form \[\frac \alpha{\beta+\sqrt \gamma }\] we have to multiply by \(\dfrac {\beta-\sqrt \gamma}{\beta-\sqrt \gamma}\) (which equals one) \[\large\frac \alpha{\beta+\sqrt \gamma }=\frac \alpha{\beta+\sqrt \gamma }\times\frac {\beta-\sqrt \gamma}{\beta-\sqrt \gamma }\\\,\\\large \qquad\qquad =\frac{\alpha(\beta-\sqrt\gamma)}{\beta^2-\gamma}\] _____________________ \(\beta+\sqrt \gamma\) and \(\beta-\sqrt \gamma \) are called irrational conjugates of each other .

OpenStudy (anonymous):

5√3/3

OpenStudy (cggurumanjunath):